# Weighted Parameter Estimators of the Generalized Extreme Value Distribution in the Presence of Missing Observations

### v1

@@ 5 Tidal Wave Surges in Eastern Canada
In this data application, we modeled extremal wave surges to predict wave surge return levels. We used hourly water level data from three locations in Atlantic Canada: Saint John (ref #: 65), Yarmouth (ref #: 365) and Port-Aux-Basques (ref #: 665) with corresponding time periods of 1941-2023 (83 years), 1965-2023 (59 years) and 1959-2023 (64 years), respectively. The number of years with missing observations for each location were 60, 45, and 56 with corresponding mean proportions of missing data in years with missing observations of $12.11\%$, $12.02\%$ and $10.67\%$, respectively. The analyzed data are publicly available and can be accessed at https://www.tides.gc.ca/. For a visual representation of the proportion of missing observations by year across the three stations and the corresponding observed maxima for each year, see Figure 2.
@@
For Saint John, both conditional weighted approaches yield much larger return level estimates, with the moment-based approach having by far the largest return level estimates. However the standard error associated with these methods has approximately doubled, which may reflect the uncertainty associated with the observed maxima. The unconditional weighted approaches yield slightly larger return level estimates than for the observed estimation approaches with similar levels of uncertainty. For the Yarmouth and Port-Aux-Basques locations, the difference between results is much more moderate, with very little difference between the likelihood-based approaches at the Port-Aux-Basques location.
@@
MLE-Complete $2.443$ $3.564$ $4.761$ $1.16\times 10^{9}$ $9.15\times 10^{12}$ $7.62\times 10^{15}$
@@
MLE-Complete $1.164$ $1.445$ $1.714$ $3.57\times 10^{10}$ $1.03\times 10^{15}$ $2.30\times 10^{18}$
@@
MLE-Complete $1.045$ $1.154$ $1.236$ $2.38\times 10^{26}$ $1.490\times 10^{36}$ $3.272\times 10^{43}$
@@ Appendix A Technical Remarks
In the general setup of Hu (1997), they assume the random variables $X_{1},...,X_{n}$ are independently drawn from densities $f_{j}$, $j=1,...,n$ which are “like” a common density function $f(\cdot;\theta)$ for $\theta\in\Omega$ where the “likeness” depends on the relevance weights $p_{nj}$, $j=1,...,n$. Similarly, in the missing data setting, we assume the observed block maxima, $X_{1},...,X_{n}$, which may not necessarily correspond to the true block maxima, are drawn from density functions $f_{1},...,f_{n}$ which are related to the true GEV distribution through the use of weighting functions. Clearly, in the unconditional weighting scenario, the weights $\hat{w}_{j}^{(1)}=\frac{n_{j}}{n_{j}+n_{j}^{\prime}}$ for $j=1,...,n$ are non-negative and although the weights do not sum to $1$, the normalized weights $\frac{\hat{w}_{j}}{\sum_{i=1}^{n}\hat{w}_{i}}$ do satisfy this constraint and result in the same likelihood estimators.
To establish consistency and asymptotic Normality of the weighted maximum likelihood estimators, Hu (1997) lists a series of approximately 20 assumptions (similar to the assumptions in the original proofs of Wald). Here, we will argue why the assumptions of Hu (1997) are valid in the missing extreme value setting by referring directly to the numbered assumptions of Hu (1997). First, we assume the individual density functions $f_{1},...,f_{n}$ have the same support as the target GEV distribution. This assumption is tied directly to the individual densities being “like” $f$ and is satisfied by the limiting distribution of $r$ largest order statistics having the same support as the limiting GEV distribution, see Galambos (1978) for proof. This assumption on the support of the density functions corresponds to condition 2.1.1 of Hu (1997). For condition 2.1.2, we assume there exists a true $\theta_{0}$ satisfying $\sum_{j=1}^{n}\hat{w}_{j}^{(1)}\mathbb{E}\log\frac{f_{j}(X_{j})}{f(X_{j};\theta_{0})}\rightarrow 0$ as $n\rightarrow\infty$ and for any $\theta$, $\sum_{j=1}^{n}\hat{w}_{j}^{(1)2}\mathbb{V}\log\frac{f_{j}(X_{j})}{f(X_{j};\theta)}\rightarrow 0$ as $n\rightarrow\infty$. These two criteria ensure that the Kullback-Leibler divergence tends to 00 as the sample size grows. As remarked in Hu (1997), since $\max_{j}\hat{w}_{j}^{(1)}\rightarrow 0$ (i.e. as more data are collected sequentially, the maximum of the block weights tends to 00), the variance condition generally holds as $\mathbb{V}\log\frac{f_{j}(X_{j})}{f(X_{j};\theta)}$ is uniformly bounded for each $\theta$. Finally, Hu (1997) makes the standard assumption of maximum likelihood estimation problems (condition 2.1.3), which assumes the parameter space $\Omega$ contains an open interval $\Theta$ where $\theta_{0}$ is an interior point. In the setting of the GEV, assuming $\xi>-0.5$, this assumption is trivially satisfied.
Assumptions 2.2.1-2.2.8. are similar to the assumptions of the Wald to establish the required lemmas needed in proving the consistency of the MLEs. Assumptions 2.2.9. and 2.2.10. also hold in the weighted GEV setting as the dominated-convergence theorem and weak law of large numbers holds in the extreme value distribution framework. Under the listed assumptions, the weighted MLE can be shown to be a weakly consistent estimator of $\theta_{0}$. The assumptions required to prove the asymptotic Normality of the MLE are again classical as they are based on the approach used by Cramer. As discussed above, assumptions 3.1.1-3.1.8 of Hu (1997) are stated in the general independent but not identically distributed case with unequal probability weights and are therefore directly applicable to our proposed maximum likelihood estimation procedure.
As commented in Section 3.4, we note that the approach taken in Hu (1997) assumes the relevance weights $p_{nj}$ are not random and are fixed by the researcher. It is not immediately clear how the classical approaches of Wald and Cramer can be adjusted to allow for random relevance weighting functions such as for the conditional weights given by $\hat{w}_{j}^{(2)}$.
@@
To establish asymptotic Normality of the Probability-Weighted Moment Estimator, Hosking et al. (1985) appealed to the proof technique of Chernoff et al. (1967) which, like the weighted maximum likelihood proof, required three conditions. We will verify that the three conditions still hold with the inclusion of the unconditional weights. For specific details on the conditions and how they relate to more general estimation problems, refer to the original work of Chernoff et al. (1967). First, condition A states that the terms inside the summation are continuously differentiable. As the inclusion of constant weighting functions does not affect differentiability or the unweighted estimator proposed in Hosking et al. (1985), this condition holds trivially. Similarly, conditions B and C state that particular summations are expressions that grow at little o rates. However, since all the unconditional weights are constant and bounded above by $1$, then these conditions will still hold as the same upper bound may be assumed. In the case where the weights are conditionally based on the observed maxima, the original conditions of Chernoff et al. (1967) may not be sufficient as the differentiability condition will depend on the conditional weighting function and the rates of growth for the summations will also depend on the inputs of the conditional weighting functions.

### v2

@@ 5 Tidal Wave Surges in Eastern Canada
[…]
@@
For Saint John, both conditional weighted approaches yield much larger return level estimates, with the moment-based approach having by far the largest return level estimates. However the standard error associated with these methods has approximately doubled, which is a reflection of the larger return level estimates and may also reflect the uncertainty associated with the observed maxima. The unconditional weighted approaches yield slightly larger return level estimates than for the observed estimation approaches with similar levels of uncertainty. For the Yarmouth and Port-Aux-Basques locations, the difference between results is much more moderate, with very little difference between the likelihood-based approaches at the Port-Aux-Basques location.
@@
MLE-Complete 2.4432.4432.443 3.5643.5643.564 4.7614.7614.761 $1.16\times 10^{9}$ $9.15\times 10^{12}$ $7.62\times 10^{15}$
@@
MLE-Complete 1.1641.1641.164 1.4451.4451.445 1.7141.7141.714 $3.57\times 10^{10}$ $1.03\times 10^{15}$ $2.30\times 10^{18}$
@@
MLE-Complete 1.0451.0451.045 1.1541.1541.154 1.2361.2361.236 $2.38\times 10^{26}$ $1.490\times 10^{36}$ $3.272\times 10^{43}$
@@ Appendix A Technical Remarks
Using the relevance maximum likelihood setup of Hu (1997), we denote the observed block maxima by $M_{n_{1}},...,M_{n_{k}}$ and assume that they are independently drawn from densities $f_{j}$, for $j=1,...,k$. As the observed maxima are not necessarily drawn directly from the GEV, given by $f(\cdot;\theta)$ for $\theta\in\Omega$, we introduce relevance weights to weight the relevance of density $f_{j}$ to the target density $f(\cdot;\theta)$. The weights are normalized as $\frac{\hat{w}_{i}}{\sum_{j=1}^{k}\hat{w}_{j}}$ for $i=1,...,k$ to ensure they are non-negative and they sum to 111. Changing from the original weights $\hat{w}_{i}$ to the normalized weights results in the same maximum likelihood estimators.
One of the main challenges in establishing the asymptotic properties of the weighted ML estimator is defining the form of the observed maxima densities, $f_{j}$, for $j=1,...,k$ in conjunction with their corresponding weighting functions. Specifically, in a standard extreme value analysis without missing observations where the block size is assumed to be large, the density of the observed maximum can be approximated by the GEV distribution. In the presence of missing observations, the observed maximum may not correspond to the true block maximum but could correspond to a lower order statistic. The way in which the data are assumed to be missing will directly impact the asymptotic distribution of the observed block maxima. For example, when the upper 10% of the series level data are always missing according to an MNAR missingness mechanism, the limiting behaviour of the observed block maximum is not immediately clear as its order will change as block size goes to infinity. The rate the order changes which depends on the missingness mechanism. Thus, the observed maxima densities must be uniquely defined for each type of missingness mechanism.
A second difficulty in establishing the asymptotic properties for the weighted maximum likelihood estimator is through the proposed weighting functions. The proposed weights are both random in that they either depend on the random proportion of observed data within a block (i.e. the unconditional weights) or directly depend on the series level data through the empirical cumulative distribution function. We note however that the relevance weighted maximum likelihood estimator described in Hu (1997) assumes the relevance weights $p_{nj}$ are not random and are fixed by the researcher. As the asymptotic proof approach is based on the original techniques of Wald and Cramer, it is not clear how the approach can be adjusted to allow for random relevance weighting functions.
@@
To establish asymptotic Normality of the Probability-Weighted Moment Estimator, Hosking et al. (1985) appealed to the proof technique of Chernoff et al. (1967) which required three conditions. We will verify that the three conditions still hold with the inclusion of the unconditional weights (assuming the number of missing/observed series level data per block are fixed). For specific details on the conditions and how they relate to more general estimation problems, refer to the original work of Chernoff et al. (1967). First, condition A states that the terms inside the summation are continuously differentiable. As the inclusion of constant weighting functions does not affect differentiability or the unweighted estimator proposed in Hosking et al. (1985), this condition holds trivially. Similarly, conditions B and C state that particular summations are expressions that grow at little o rates. However, since all the unconditional weights are constant and bounded above by 111, then these conditions will still hold as the same upper bound may be assumed. In the case where the weights are conditionally based on the observed maxima, the original conditions of Chernoff et al. (1967) may not be sufficient as the differentiability condition will depend on the conditional weighting function and the rates of growth for the summations will also depend on the inputs of the conditional weighting functions.